# Tradeoffs Between Alignment and Helpfulness in Language Models

### v2

@@ Abstract
Language model alignment has become an important component of AI safety, allowing safe interactions between humans and language models, by enhancing desired behaviors and inhibiting undesired ones. It is often done by tuning the model or inserting preset aligning prompts. Recently, representation engineering, a method which alters the model’s behavior via changing its representations post-training, was shown to be effective in aligning LLMs (Zou et al., 2023a). Representation engineering yields gains in alignment oriented tasks such as resistance to adversarial attacks and reduction of social biases, but was also shown to cause a decrease in the ability of the model to perform basic tasks. In this paper we study the tradeoff between the increase in alignment and decrease in helpfulness of the model. We propose a theoretical framework which provides bounds for these two quantities, and demonstrate their relevance empirically. Interestingly, we find that while the helpfulness generally decreases, it does so quadratically with the norm of the representation engineering vector, while the alignment increases linearly with it, indicating a regime in which it is efficient to use representation engineering. We validate our findings empirically, and chart the boundaries to the usefulness of representation engineering for alignment.
@@ 1 Introduction
Advancements in large language model (LLM) development over the last few years have given LLMs a variety of abilities that allow them to serve as general purpose assistants in a wide range of tasks, such as broad-scoped question answering, writing assistance, teaching, and more  (Radford et al., 2019; Devlin et al., 2019; Brown et al., 2020; Schulman et al., 2023; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023; Park et al., 2023a). The vast use of LLMs for such purposes has raised concerns due to the harm they can cause their users, such as feeding fake information (Lin et al., 2022; Weidinger et al., 2022), behaving offensively and feeding social biases (Hutchinson et al., 2020; Venkit et al., 2022; Weidinger et al., 2022), or encouraging problematic behaviors by users Roose (2023); Atillah (2023). Alignment is often the term given for the process of removing these undesired behaviors (Yudkowsky, 2001; Taylor et al., 2016; Amodei et al., 2016; Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Pan et al., 2022; Ngo, 2022).
@@
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021); Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b) and the work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
@@
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, in order to steer from undesired to desired behaviors. With it, they demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023b); Turner et al. (2023) to improve alignment and reduce toxicity.
@@
There are also known limitations to editing representations. The work of Elazar et al. (2021) empirically demonstrates how projecting out supervised linear probe directions via iterative nullspace projection can reduce performance on selected tasks. Understanding the tradeoff between model helpfulness and alignment is important for designing safe LLM systems that incorporate representation engineering. In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
In sections 3 and 4, we set up our theoretical framework and theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective, allowing for inference time alignment while maintaining the model’s helpfull capabilities. See figure 1 for an illustration of this intuition.
In section 5 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by 1. Then, we calculate the average helpfulness over datasets of questions from different domains as a function of injected vector norms for the same behaviors and find that the decay with increased vector norm described in theorem 2 is manifested. Together, these results correspond to the intuitive illustration in figure 1 .
2 Related Work on Reperesentation Engineering
The so-called method of representation engineering was proposed by Zou et al. (2023a) as an alternative method for LLM alignment. Representation editing for LLM alignment purposes has been used in several other notable works: Jorgensen et al. (2023) show that using this method with the mean of the representation differences as steering vectors can more efficiently reduce toxicity, and Leong et al. (2023) also use this method for reducing toxicity. Liu et al. (2023b) combine this method with human feedback to improve model’s behaviors to human preferences. A similar method had been proposed by Turner et al. (2023) which subtracts layer activations in order to steer a model in latent space towards more desirable behaviors. These works propose steering directions by subtracting hidden states of prompts that display opposite behaviors. Thus, as they show experimentally, when added to the representation of a new query, its representation is steered in latent space towards a direction that leads to a more desirable behavior from the model output.
The method of representation editing is more veteran than the above alignment related efforts; it was previously used for style transfer (Shen et al., 2017), knowledge editing (Hernandez et al., 2023), controlled text-generation (Dathathri et al., 2019), truthfulness enhancement (Li et al., ), replacement of in-context learning prompts with in-context learning vectors (Liu et al., 2023a), and more. Theoretical works on the subject show that in context learning is equivalent to inserting a query with a task vector Hendel et al. (2023) and the vectors that connect representations of token pairs are parallel when the semantic difference between the pairs is similar (the linear representation hypothesis) Park et al. (2023b).
3 Necessary Preliminaries
@@
Where $r^{(L)}_{\theta}(s)$ is the final hidden layer’s representation of the prompt $s$ and $U$ is a matrix from the hidden state to a vocabulary of tokes. This is an accurate parametrization for state-of-the-art LLMs.
@@
Note that here $l<L$, as used in Zou et al. (2023a). Additionally, we follow existing methods for representation engineering and provide a uniform norm for all the injected vectors $|r_{e}^{(l)}|=|r_{e}|$ (where the coefficient $r_{e}$ can be positive or negative) or $|r_{e}^{(l)}|=0$ for layers that are not injected.
We use the behavior expectation definition of alignment as in Wolf et al. (2023). Here we will use a binary behavior scoring function:
@@
Let $B:\Sigma\rightarrow\{-1,+1\}$ be a binary behavior scoring function, the behavior of a prompted model $\mathbb{P}(\cdot|q)$ is defined as:
$B[\mathbb{P}(\cdot|q)]=\mathbb{E}_{a\sim\mathbb{P}(\cdot|q)}[B(a)]=$ (3)
$=\sum_{a_{+}\in aligned}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in misaligned}\mathbb{P}_{\theta}(a_{-}|q)$ (4)
Notice that the behavior expectation is in the range $[-1,+1]$. In 1 we will prove that representation engineering is an effective method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputing the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$.
@@ 4 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 2 that as the norm is increased, the helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. If the change to the final hidden layer due to $r_{e}$ classifies the representations of aligned and misaligned answers with respect to $B$, with margin $\Delta$, the behavior expectation of the model conditioned on the query $q$ satisfies:
@@
As can be seen in the mathematical expression and in figure 2 for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. For proof see appendix section B.
Figure 2: Plot of the upper bound on behavior expectation in theorem 1
Note that while the theorem takes into account the change to the final hidden layer representation, this change incorporates the changes due to the steering vectors from all previous layers, $R_{e}=\{r_{e}^{(1)},...,r_{e}^{(L-1)}\}$.
Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in E, to yield similar results.
Now, we shall bound from above the helpfulness of the model as a function of representation engineering.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. The helpfulness of the model on the query is bounded with probability $1-\frac{2}{|V|}$ by:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)\leq\frac{P_{0}}{P_{0}+(1-P_{0})\cdot\alpha(1+\frac{\beta^{2}}{2}r_{e}^{2})}$ (7)
Where $P_{0}=\mathbb{P}_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $|V|$ is the vocabulary size and $\alpha,\beta>0$ that depend on the query.
The above bound is illustrated in figure 3 for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$.
Importantly, this demonstrates that for large $r_{e}$, the helpfulness decays to zero, hence representation engineering significantly harms the model’s overall performance, while for small $r_{e}$, it can initially decrease more slowly (parabolically) around $r_{e}=0$, hence the model’s performance is relatively unharmed. For the second statement to be feasible, the true helpfulness and the bound need to be close when no representation engineering is performed, such that the parabolic decrease of the upper bound forces the helpfulness to decrease as well around $r_{e}=0$. Generally, the difference between the two at $r_{e}=0$ is bounded by $1-P_{0}$, such that for queries with high probability of being answered correctly without representation engineering, i.e. $P_{0}\approx 1$, the true helpfulness and the bound will be close, guaranteeing the parabolic bound to be meaningful.
@@
The parameter $\beta$ is a measure for the rate/curvature of the quadratic decay, as it is the coefficient multiplying $r_{e}^{2}$. It is the minimum between two weighted sums of half normal distributions, whose variance $\sigma^{2}$ depends on the dataset. In section A.3, we empirically find that $\sigma\approx 0.5$, such that the mean of each weighted sum is $\approx 0.4$ (mean of a half normal distribution), therefor $\beta$ is not too small and the decay should be felt at coefficients $r_{e}$ of order of size $1$.
Figure 3: Plot of helpfulness bound with given parameters of $P_{0}$, $\alpha$ and $\beta$.
@@
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq\textit{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ (8)
 $\textit{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}(1-\frac{2}{|V|})\frac{1}{1+\alpha(1+\frac{\beta^{2}}{2}r_{e}^{2})}+\frac{2}{|V|}$ (9)
For $\alpha,\beta>0$ that depend on the dataset.
So for datasets that have high helpfulness expectation value (about $0.9$ or higher), this bound shows helpfulness decay quadratically.
@@
Under assumption 3, the expected value of the helpfulness on a dataset of queries, $\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]$ is asymptotically bounded from above by $\frac{1}{N}$ as $|r_{e}|\rightarrow\infty$. Where $N$ is the number of possible answers for each query.

### v3

@@ Abstract
Language model alignment has become an important component of AI safety, allowing safe interactions between humans and language models, by enhancing desired behaviors and inhibiting undesired ones. It is often done by tuning the model or inserting preset aligning prompts. Recently, representation engineering, a method which alters the model’s behavior via changing its representations post-training, was shown to be effective in aligning LLMs (Zou et al., 2023a). Representation engineering yields gains in alignment oriented tasks such as resistance to adversarial attacks and reduction of social biases, but was also shown to cause a decrease in the ability of the model to perform basic tasks. In this paper we study the tradeoff between the increase in alignment and decrease in helpfulness of the model. We propose a theoretical framework which provides bounds for these two quantities, and demonstrate their relevance empirically. First, we find that under the conditions of our framework, alignment can be guaranteed with representation engineering, and at the same time that helpfulness is harmed in the process. Second, we show that helpfulness is harmed quadratically with the norm of the representation engineering vector, while the alignment increases linearly with it, indicating a regime in which it is efficient to use representation engineering. We validate our findings empirically, and chart the boundaries to the usefulness of representation engineering for alignment. Our code is available at https://github.com/dorin133/REPE_alignment_helpfulness_tradeoff.
@@ 1 Introduction
Advancements in large language model (LLM) development over the last few years have given LLMs a variety of abilities that allow them to serve as general purpose assistants in a wide range of tasks, such as broad-scoped question answering, writing assistance, teaching, and more  (Radford et al., 2019; Devlin et al., 2019; Brown et al., 2020; Schulman et al., 2023; OpenAI, 2023; Bubeck et al., 2023; Nori et al., 2023; West, 2023; Park et al., 2023a). The vast use of LLMs for such purposes has raised concerns due to the harm they can cause their users, such as serving fake information (Lin et al., 2022; Weidinger et al., 2022), behaving offensively, feeding social biases (Hutchinson et al., 2020; Venkit et al., 2022; Weidinger et al., 2022), or encouraging problematic behaviors by users Roose (2023); Atillah (2023). Alignment is often the term given for the process of removing these undesired behaviors (Yudkowsky, 2001; Taylor et al., 2016; Amodei et al., 2016; Shalev-Shwartz et al., 2020; Hendrycks et al., 2021; Pan et al., 2022; Ngo, 2022).
@@
There are several different approaches to performing alignment in LLMs, such as including aligning prompts Askell et al. (2021); Rae et al. (2021) which was shown to improve alignment and decrease toxicity in LLMs, and the procedure of reinforcement learning from human feedback (RLHF) which trains language models to be helpful and harmless (Bai et al., 2022). Though effective to an extent, these approaches are still dangerously frail, as several works have shown that adversarial prompts can trigger negative behaviors in LLMs  Wallace et al. (2019); Yu & Sagae (2021); Xu et al. (2021); Subhash (2023); Zou et al. (2023b). The work of Wolf et al. (2023) provides a theoretical framework which shows that frozen LLMs can be misaligned with sufficiently long prompts.
@@
Recently, a new method for alignment has been proposed by Zou et al. (2023a), which controls the model at the internal representation level by adding tailored vectors to the hidden layers’ representations. This is done by extracting directions in the model’s latent space that connect contrasting behaviors, and then injecting vectors at inference time in order to steer away from undesired behaviors and towards desired ones. Typically, the vectors are all prepared with norm $1$, and they are multiplied by a coefficient to tune the strength of the steering, but there is a trade-off – when the parameter is too high performance tends to degrade. Zou et al. (2023a) demonstrated experimentally that the procedure can significantly improve alignment, e.g., in resistance to adversarial attacks, with reduction from 50% success of adversarial attacks to less than 15%, and truthfulness enhancement, with a relative increase of over 50%, though at the cost of somewhat reducing the helpfulness of the model. Similar methods have also been used by Jorgensen et al. (2023); Leong et al. (2023); Liu et al. (2023); Turner et al. (2023) to improve alignment and reduce toxicity.
Since then, there has been an increasing body of work using this method. Wang et al. (2024b) use extracted safety vectors for inference time alignment for harmlessness, reducing jailbreaking success rate from over 30% with prompting and over 10% in supervised fine tuning to below one percent. Wang et al. (2024a) uses a method of editting model parameters that maximize the difference between toxic and untoxic responses to detoxify it. Wei et al. (2024) find sparse regions in parameter space that affect alignment brittleness, to be removed for better alignment. Marks et al. (2024) interpret causal graphs in language models and edit them to improve behaviors. van der Weij et al. (2024) extend activation steering to multiple behaviors. To improve low rank finetuning, Wu et al. (2024) utilize a procedure of tuning representations directly to substantially reduce the trainable parameters of finetuning compared to LoRA. Xu et al. (2024); Li et al. (2024) use concept activation vectors to jailbreak, they also observe that concepts that activate different behaviors are linearly separable. Zhang et al. (2024) remove hallucinations by editting truthfulness concepts.
There are also known limitations to editing representations. Limitations of model editing methods for social debiasing are studied in Yan et al. (2024), and the work of Elazar et al. (2021) empirically demonstrates how projecting out supervised linear probe directions via iterative nullspace projection can reduce performance on selected tasks. Theoretical works on the subject show that in context learning is equivalent to inserting a query with a task vector Hendel et al. (2023) and the vectors that connect representations of token pairs are parallel when the semantic difference between the pairs is similar (the linear representation hypothesis) Park et al. (2023b).
@@
Understanding the tradeoff between model helpfulness and alignment is important for designing safe yet useful LLM systems. In this work we aim to shed light on the benefits and limitations of representation engineering for LLM alignment, i.e., how much does alignment improve with this method and what is the cost in terms of the model’s abilities. We approach this question theoretically at first, and then provide empirical evidence for the validity of our theory.
@@
In sections 2 and 3, we set up our theoretical framework and present our theoretical results respectively. We find that representation engineering increases alignment linearly with the representation engineering vector norm (theorem 1), while the helpfulness of the model, defined as the probability of answering a query correctly, decreases quadratically with the representation engineering vector norm (theorem 2). Consequently, alignment can be guaranteed with large enough representation injections (corollary 1), though at the cost of significantly reducing the model’s helpfulness, e.g.multiple choice question answering reduced to random guessing (corollary 3). Conversely, this means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective, allowing for inference time alignment while maintaining the model’s helpful capabilities. See figure 1 for an illustration of this intuition.
In section 4 we explore the validity of our assumptions and results in an experimental setting: We calculate alignment, as defined by the theoretical framework, as a function of representation engineered vector norms corresponding to the desired behaviors and find that it increases as predicted by theorem 1. This is done by aligning with representation engineering an unaligned (pretrained) model with respect to desired behaviors (“harmless", “not-racist"), and misaligning an aligned (RLHF) model to undesired behaviors (“harmful", “racist"). Then, we calculate the helpfulness of the model, quantified by its question answering abilities over different knowledge domains, with the same aligning vectors, and find that the decay with increased vector norm described in theorem 2 is manifested. Furthermore, injecting large norms of these vectors leads to significant reduction of helpfulness, such as performance on multiple-choice questions that is equivalent to random guessing (corollary 3). Together, these results correspond to the intuitive illustration in figure 1 .
2 Preliminaries
@@
Where $r^{(L)}_{\theta}(s)$ is the final hidden layer’s representation of the prompt $s$ and $U$ is a matrix from the hidden state to a vocabulary of tokens. This is an accurate parametrization for state-of-the-art LLMs.
@@
Note that here $l<L$, as used in Zou et al. (2023a). Additionally, we follow existing methods for representation engineering and provide a uniform norm for all the injected vectors $|r_{e}^{(l)}|=|r_{e}|$. The vectors are initially prepared with norm $1$, and when injected to the model, they are multiplied by the coefficient $r_{e}$ which can be positive or negative, to tune the steering strength and direction. For layers that are not injected, $|r_{e}^{(l)}|=0$.
To quantify alignment, we use the behavior expectation definition of alignment as in Wolf et al. (2023). We will use a binary behavior scoring function:
@@
Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a binary behavior scoring function, the behavior of a prompted model $\mathbb{P}(\cdot|q)$ is defined as:
$B[\mathbb{P}(\cdot|q)]=\mathbb{E}_{a\sim\mathbb{P}(\cdot|q)}[B(a)]=\sum_{a_{+}\in aligned}\mathbb{P}_{\theta}(a_{+}|q)-\sum_{a_{-}\in misaligned}\mathbb{P}_{\theta}(a_{-}|q)$ (3)
Notice that while $B$ is a binary function, the behavior expectation is in the range $[-1,+1]$. A binary behavior scoring function reflects cases where an aligned response is required and unaligned responses must be filtered. In theorem 1 we will prove that representation engineering is an effective method by lower bounding the behavior expectation. Notice that high probability of outputting a positive/negative response gives a positive/negative contribution to the behavior expectation, thus the sign and absolute value of behavior expectation is a good measure for the alignment of a model.
@@
Where $P_{\theta}(a_{correct}|q)$ is the model’s probability of outputting the correct answer $a$ to the query $q$. By this definition, the helpfulness is in the range $[0,1]$. The motivation behind this definition is to quantify the general capabilities of the model when engineered representations are injected into it.
Ideally, a model that interacts with a user should be both aligned and helpful, meaning its response is appropriate w.r.t. a desired behavior (i.e., positive behavior expectation) and also useful (i.e., high probability of giving a correct answer to the query).
Lastly, note the use of the model’s next token prediction to determine alignment and helpfulness. Initially this may seem to restrict the generality of our results, but note that for alignment, it is often the first token that determines whether an answer is aligned or not ("Sorry, as an AI I cannot…" vs "Sure! Here’s…"). For helpfulness this point is more subtle, as some tasks are more intricate, however, as demonstrated by our experiments, there are numerous tasks that can be reduced to next token prediction.
@@ 3 Main Results
We will show that representation engineering improves alignment, but harms helpfulness. Theorem 1 shows that behavior expectation is bounded from below by a hyperbolic tangent function, such that it approaches $+1$ for increasing size of injected vectors and increases linearly within a bounded range. This in principle allows to sample an aligned response for any adversarial attack (corollary 1), demonstrating the power of representation engineering as an alignment technique. Theorem 2 shows that the helpfulness is maximized in the vicinity of norm zero injected vectors (i.e., no representation engineering) and in corollary 3 that as the norm is increased, the helpfulness decays to random guessing. The assumptions used to prove the theorems are presented formally in appendix A.
The following statement quantifies how alignment is improved by representation engineering. It assumes the change in all layers amounts to a good classifier of positive and negative behavior tokens in the last layer w.r.t.the query, as depicted in figure 2a. This condition was chosen due to the popular choice in representation engineering to use injected representations $\{r_{e}^{(l)}\}$, that are themselves classifiers for positive and negative representations on the intermediate layers. This is because they are learned from contrasting positive and negative behavior representations for different queries, such as mean centering, $r_{e}^{(l)}=\mathbb{E}_{good,bad}[r_{good}^{(l)}-r_{bad}^{(l)}]$, or PCA, $r_{e}^{(l)}=\text{argmax}_{v:||v||=1}[\mathbb{E}_{good,bad}|\langle v,r_{good}^{(l)}-r_{bad}^{(l)}\rangle|^{2}]$ such that they form linear classifiers for the intermediate layers due to the positive/negative inner product with positive/negative answer representations. Notably, in Xu et al. (2024) it is shown empirically that such concept classes in latent space are linearly separable. We discuss this assumption further in A and provide empirical evidence. Furthermore, the classification condition can be softened to an imperfect classifier, as discussed in appendix A and shown in appendix in F, to yield similar results.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. Let $B:\Sigma^{*}\rightarrow\{-1,+1\}$ be a behavior scoring function. The injections to all layers amounts to a change in the final hidden layer representation that is $q$ dependent, denoted by the vector $\delta r^{(L)}_{e}(q)$. Let $\Delta$ be the margin of classification over the model’s representations for aligned and misaligned answers with respect to $B$, when using $\delta r^{(L)}_{e}(q)$ as a classifier. Then, the behavior expectation of the model conditioned on the query $q$ satisfies:
@@
As can be seen in the mathematical expression and in figure 2b for $B_{0}=-0.5$, this lower bound is a shifted hyperbolic tangent function w.r.t $r_{e}$. At $r_{e}=0$ the bound gives $B_{0}$, which is the unaltered model’s behavior. As $r_{e}$ is increased, the bound approaches $+1$, meaning the behavior asymptotically approaches $+1$. We see that for $B_{0}$ that is not too close $-1$, the increase in behavior expectation is linear due to the hyperbolic tangent’s nature, while if it is very close to $-1$, $r_{e}$ is to be increased before seeing the linear effect. Thus for behaviors on which the model is negative but also has a small tendency for positive answers, the linear effect should be felt near $r_{e}=0$. In section 4, we present our numerical estimation $\Delta\lambda$ in the range $0.1-3$, both based on the linear classifier condition and direct alignment measurement. For proof see appendix section B.
Figure 2: (a) The change to the last hidden layer due to vector injections from previous layers classifies positive and negative answer representations. (b) Plot of the upper bound on behavior expectation in theorem 1
Note that during decoding, the change in all layers of the model $R_{e}=\{r_{e}^{(1)},...,r_{e}^{(L-1)}\}$, amounts to a change to the final hidden layer’s representation, $r_{\theta,r_{e}}^{(L)}-r_{\theta}^{(L)}$, where $r_{\theta,r_{e}}^{(L)}$ is the final hidden layer representation that incorporates all the previous hidden layer changes and $r_{\theta}^{(L)}$ is the original representation. The complexity of the multi-layer editing is incorporated into the types of changes observed in the final hidden layer.
We note that in contrast to Wolf et al. (2023), that has a framework centralized on using prompts to misalign frozen models, i.e.models whose weights and representations are not changed after training, here the model is not frozen due to representation engineering, and accordingly a different result is obtained on guaranteeing an aligned response - for any adversarial attack, it is possible to use large enough norms with representation engineering to produce an aligned response if the learned injected representations accumulate to a good classifier of positive and negative answer representations in the final layer.
Corollary 1.
Let $\epsilon>0$, $\mathbb{P}_{\theta}$ a language model and $q$ a prompt that induces negative behavior $B[\mathbb{P}_{\theta}(\cdot|q)]<\gamma<0$ without representation engineering. Under the conditions of theorem 1, using an injected vector norm of $r_{e}>\frac{1}{\Delta\lambda}(arctanh(1-\epsilon)-arctanh(\gamma))$ leads to behavior expectation $B[\mathbb{P}_{\theta,r_{e}}(\cdot|q)]>1-\epsilon$.
Now, we shall bound from above the helpfulness of the model as a function of representation engineering. Intuitively, editting the model’s representation in a specific direction adds random noise to other latent concepts of the model, causing a degradation in its other capabilities. The following theorem formally states this.
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a model prompted with query $q$ and injected with representations of coefficient $r_{e}$. If the directionality of the last hidden layer representation change due to the injections in all layers, distributes normally with variance $\sigma^{2}$ w.r.t. the correct and incorrect answer representations, the helpfulness of the model on the query is bounded with probability $1-\frac{2}{|V|}$ by:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)\leq\frac{P_{0}}{P_{0}+(1-P_{0})\cdot\alpha(1+\frac{\lambda^{2}\sigma^{2}\beta^{2}}{2}r_{e}^{2})}$ (6)
Where $P_{0}=\mathbb{P}_{\theta,r_{e}=0}(\cdot|q)$ is the probability of answering correctly without representation engineering, $|V|$ is the vocabulary size and $\alpha,\beta>0$ that depend on the query. $\lambda$ is a model dependent coefficient relating between $r_{e}$ and the corresponding final hidden state norm.
The proof is presented in appendix C. The above bound is illustrated in figure 3 for different values of $\beta$. As can be seen, around $r_{e}=0$, the bound is parabolic, i.e.  the decrease is proportional to $-r_{e}^{2}$, this can be obtained by expanding the bound near $r_{e}=0$. On the other hand, for large $r_{e}$, we see a decay to zero at a rate proportional to $r_{e}^{-2}$, this can be obtained by expanding the bound for large $r_{e}$.
Importantly, this demonstrates that for large $r_{e}$, the helpfulness decays to zero, hence representation engineering significantly harms the model’s overall performance, while for small $r_{e}$, it can initially decrease more slowly (parabolically) around $r_{e}=0$, hence the model’s performance is relatively unharmed. For the second statement to be feasible, the true helpfulness and the bound need to be close when no representation engineering is performed. The difference between the two at $r_{e}=0$ is bounded by $1-P_{0}$, such that for queries with high probability of being answered correctly without representation engineering, i.e. $P_{0}\approx 1$, the true helpfulness and the bound will be close, guaranteeing the parabolic bound to be meaningful.
@@
The product of parameters $\lambda\sigma\beta$ are a measure for the rate/curvature of the quadratic decay, as they are the coefficient multiplying $r_{e}^{2}$. $\lambda$ is the same scaling parameter from theorem 1, $\sigma$ represents the standard deviation of random noise added to the logits due to representation engineering (depicted in figure 3a and formally defined in A), and $\beta$ is the minimum between two weighted sums of half normal distributions with parameter $\sigma^{\prime}=1$. In section 4, we present our empirical estimation $\lambda\sigma\beta$ in the range $0.1-0.66$, both based on the normal logit noise condition and direct helpfulness measurement. Hence the decay should be felt at coefficients $r_{e}$ of order of size $1$.
Figure 3: (a) Directionality of change to last hidden layer due to representation engineering distributes normally with variance $\sigma^{2}$ w.r.t. correct and incorrect answer representations. (b) Plot of helpfulness bound with given parameters of $P_{0}$, $\alpha$ and $\lambda\sigma\beta$.
@@
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq(1-\frac{1}{|V|})\frac{1}{1+\alpha^{\prime}(1+\frac{\sigma^{2}\lambda^{2}\beta^{\prime 2}}{2}r_{e}^{2})}+\frac{1}{|V|}$ (7)
For $\alpha^{\prime},\beta^{\prime}>0$ that depend on the dataset.
So, for datasets that have high helpfulness expectation value (about $0.9$ or higher), this bound shows helpfulness decay quadratically.
@@
Under the conditions of theorem 2, the expected value of the helpfulness on a dataset of queries, $\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]$ is asymptotically bounded from above by $\frac{1}{N}$ as $|r_{e}|\rightarrow\infty$. Where $N$ is the number of possible answers for each query.